OpenStudy (anonymous):

square painting has an area of 1040 in2. To the nearest tenth, what is the length of 1 side of the painting?

OpenStudy (trojanpoem):

The area of square = (sidelength)^2 1040 = L^2 L = rootof(1040) L= 32.2

OpenStudy (anonymous):

Alex wants to put a fence around a square garden. The garden has an area of 784 ft2. What is the perimeter of the garden?